How do I get this number?
I am a three-digit even number.
I am made up of three different numerals.
I am divisible by eleven.
What number am I?

Answer by Edwin McCravy(20056)

There are 28 different possibilities:

1.  132 is a 3 digit even number made up of 3 different numerals, 1, 3, and 2 and is divisible by 11 since 132 = 11*12
2.  154 is a 3 digit even number made up of 3 different numerals, 1, 5, and 4 and is divisible by 11 since 154 = 11*14
3.  176 is a 3 digit even number made up of 3 different numerals, 1, 7, and 6 and is divisible by 11 since 176 = 11*16
4.  198 is a 3 digit even number made up of 3 different numerals, 1, 9, and 8 and is divisible by 11 since 198 = 11*18
5.  264 is a 3 digit even number made up of 3 different numerals, 2, 6, and 4 and is divisible by 11 since 264 = 11*24
6.  286 is a 3 digit even number made up of 3 different numerals, 2, 8, and 6 and is divisible by 11 since 286 = 11*26
7.  308 is a 3 digit even number made up of 3 different numerals, 3, 0, and 8 and is divisible by 11 since 308 = 11*28
8.  352 is a 3 digit even number made up of 3 different numerals, 3, 5, and 2 and is divisible by 11 since 352 = 11*32
9.  374 is a 3 digit even number made up of 3 different numerals, 3, 7, and 4 and is divisible by 11 since 374 = 11*34
10.  396 is a 3 digit even number made up of 3 different numerals, 3, 9, and 6 and is divisible by 11 since 396 = 11*36
11.  418 is a 3 digit even number made up of 3 different numerals, 4, 1, and 8 and is divisible by 11 since 418 = 11*38
12.  462 is a 3 digit even number made up of 3 different numerals, 4, 6, and 2 and is divisible by 11 since 462 = 11*42
13.  506 is a 3 digit even number made up of 3 different numerals, 5, 0, and 6 and is divisible by 11 since 506 = 11*46
14.  528 is a 3 digit even number made up of 3 different numerals, 5, 2, and 8 and is divisible by 11 since 528 = 11*48
15.  572 is a 3 digit even number made up of 3 different numerals, 5, 7, and 2 and is divisible by 11 since 572 = 11*52
16.  594 is a 3 digit even number made up of 3 different numerals, 5, 9, and 4 and is divisible by 11 since 594 = 11*54
17.  638 is a 3 digit even number made up of 3 different numerals, 6, 3, and 8 and is divisible by 11 since 638 = 11*58
18.  682 is a 3 digit even number made up of 3 different numerals, 6, 8, and 2 and is divisible by 11 since 682 = 11*62
19.  704 is a 3 digit even number made up of 3 different numerals, 7, 0, and 4 and is divisible by 11 since 704 = 11*64
20.  726 is a 3 digit even number made up of 3 different numerals, 7, 2, and 6 and is divisible by 11 since 726 = 11*66
21.  748 is a 3 digit even number made up of 3 different numerals, 7, 4, and 8 and is divisible by 11 since 748 = 11*68
22.  792 is a 3 digit even number made up of 3 different numerals, 7, 9, and 2 and is divisible by 11 since 792 = 11*72
23.  814 is a 3 digit even number made up of 3 different numerals, 8, 1, and 4 and is divisible by 11 since 814 = 11*74
24.  836 is a 3 digit even number made up of 3 different numerals, 8, 3, and 6 and is divisible by 11 since 836 = 11*76
25.  902 is a 3 digit even number made up of 3 different numerals, 9, 0, and 2 and is divisible by 11 since 902 = 11*82
26.  924 is a 3 digit even number made up of 3 different numerals, 9, 2, and 4 and is divisible by 11 since 924 = 11*84
27.  946 is a 3 digit even number made up of 3 different numerals, 9, 4, and 6 and is divisible by 11 since 946 = 11*86
28.  968 is a 3 digit even number made up of 3 different numerals, 9, 6, and 8 and is divisible by 11 since 968 = 11*88


Edwin